OpenStudy (anonymous):

Use the distributive property to simplify the expression below: 5t − 6(1 + t)

OpenStudy (goten77):

let me type this out

OpenStudy (andriod09):

\[5t-6(1+t)=0\]\[5t-6+6t=0\]\[11t-6=0\]\[11t=-6\]\[t=-0.5454545454545454\]\[\uparrow \infty\]

OpenStudy (goten77):

aah typed it wrong 1st time

OpenStudy (goten77):

@android there is no = sign so you cant solver for t

OpenStudy (andriod09):

yes, but either way, it would be equal to 0

OpenStudy (anonymous):

Use the distributive property to simplify the expression below: 5t − 6(1 + t) −t − 6 −t + 6 t + 6 6t − 6

OpenStudy (andriod09):

by me saying 1t+5000 that means 1t+5000=0

OpenStudy (goten77):

i gtg atm but im tbh u cant just assume it always = 0 and i can show exp later

OpenStudy (goten77):

.. android still did #1 wrong.... heres how u do it \[5t - 6(1+t) = 5t -6 - 6t = -t -6\] the negative has to be used.... i hope msmeloidy reads this so she isnt taught wrong

OpenStudy (andriod09):

A)ik what i did, B)it wasn't wrong, C)i didn't forget a (-)

OpenStudy (goten77):

if yours was right then y does it not match up w/ any answers??? gg there4 ur argument is invalid

hartnn (hartnn):

@andriod09 since there is no = sign, you cannot solve for x, you can only simplify here as correctly done by @Goten77